OpenStudy (anonymous):

Medal to anyone that helps! Divide. x^2-4 over x-8 divided by (x-2) over x-3

OpenStudy (anonymous):

@jcpd910 please help please

OpenStudy (anonymous):

I think your equation is: \[\frac{x ^{2}-4 }{ x-8 } \div \frac{ x-2 }{ x-3 }\] Is that correct?

OpenStudy (anonymous):

\[\frac{ (x + 2)(x -3)}{ x-8 }\]